OpenStudy (anonymous):

2x-4=3y+6 please write this equation in standard form

jimthompson5910 (jim_thompson5910):

2x-4=3y+6 2x=3y+6+4 2x-3y=6+4 2x-3y=10

OpenStudy (anonymous):

(2x-4-5)/3=y

OpenStudy (anonymous):

Could you please ask a specific question, it looks like you're having us do your homework

OpenStudy (anonymous):

everytime someone answers my questions correctly they get a metal. its a win win situation

OpenStudy (anonymous):

Not when you fail your test!